# Can Global XAI Methods Reveal Injected Bias in LLMs? SHAP vs Rule Extraction vs RuleSHAP

### v1

@@ 4 RuleSHAP
SHAP [10] is a model-agnostic XAI framework based on cooperative game theory. Given a model $f$ and an instance $u$, SHAP assigns each feature $u_{i}$ a Shapley value $\rho_{i}$, representing its average marginal contribution to the prediction $f(u)$ across all feature coalitions. This is estimated by replacing features with predefined background values (which are not necessarily 00) to measure the impact of their removal. For global attributions, SHAP averages local attributions over multiple instances [35], providing an overall measure of feature importance. These values are often visualized in summary plots like bar charts or beeswarm plots. Since computing exact Shapley values is costly and assumes feature independence, practical approximations like KernelSHAP, which employs weighted linear regression to approximate contributions, or permutation-based methods that limit feature permutations to second-order interactions, i.e., the permutation SHAP explainer (used in our experiments). To mitigate independence assumption errors, correlated features can be grouped before computation.
RuleFit [11] is a global XAI method that extracts the rules a model $f$ follows to produce outputs $\mathbf{v}$ from inputs $\mathbf{u}$. It does this by leveraging features within $\mathbf{u}$, where each input vector consists of multiple features $c_{1},c_{2},\dots,c_{m}$. RuleFit combines gradient boosting [12] with sparse linear modelling to build interpretable models. Gradient boosting iteratively constructs decision trees, where each split minimizes impurity or maximizes loss reduction. A decision path from root to leaf defines a binary rule, e.g., $(c_{i}\leq\theta_{1})\land(c_{j}>\theta_{2})\land\dots$, where $\theta_{1},\theta_{2},\dots$ are threshold values. After extracting rules, RuleFit forms a binary design matrix $\mathbf{X}$, with rows as instances and columns as rules. Each entry $X_{k,n}$ is 1 if the $k$-th rule is activated for the $n$-th instance, otherwise 0. $\mathbf{X}$ has dimensions $|R|\times|\mathbf{u}|$ (where $R$ is the set of extracted rules), capturing the presence of each rule across all instances. Then, given the coefficient vector $\mathbf{w}$ and an hyper-parameter $\alpha$ to control for sparsity, RuleFit applies a LASSO regression to select a sparse set of predictive rules by solving: $\min_{\mathbf{w}}\ \frac{1}{2}\|\mathbf{v}-\mathbf{X}\mathbf{w}\|_{2}^{2}+\alpha\|\mathbf{w}\|_{1}.$
@@
Step 1: Global Shapley Values Aggregation. We first compute SHAP values for all available data points using one of the SHAP explainers available in the official repository. Specifically, for each feature $c_{i}$ and instance $\mathbf{u}_{k}$, we obtain the Shapley value $\rho_{(i,k)}$, using the minimum feature values from $\mathcal{T}$ as the background since all features are, by construction, ordinal. Because we work with abstracted features, SHAP cannot directly estimate the average marginal contribution $\rho_{(i,k)}$, by simple feature “removal”, sending a new $\widehat{\mathbf{u}}_{k}$-vector to the LLM for evaluation.
Instead, following the workaround from Section 3, our strategy is to guarantee sufficient redundancy in $\mathcal{T}$ so that we can find multiple points $j$ in $\mathcal{T}$ for which $\|\widehat{\mathbf{u}}_{k}-\mathbf{u}_{j}\|_{2}$ is minimal ($\approx 0$) while matching the background values considered by SHAP perturbations. SHAP can then randomly select one of these points for contribution estimation, mimicking the non-determinism of LLMs and enabling a statistical assessment of how input abstractions correlate with bias-related properties. Specifically, we aggregate the $\rho_{(i,k)}$ values across all instances by calculating the mean of the absolute Shapley values, adding the standard deviation to capture the upper bound of each feature’s importance:
$\rho_{i}^{\text{agg}}=\frac{1}{N}\sum_{k=1}^{N}|\rho_{(i,k)}|+\text{std}\left(\{|\rho_{(i,k)}|\}_{k=1}^{N}\right)$
where $N$ is the total number of data points $\mathbf{u}$. The aggregated Shapley values $\rho_{i}^{\text{agg}}$ are then normalized to the interval $(0,1]$ by: $\bar{\rho_{i}}=\rho_{i}^{\text{agg}}/\sum_{j=1}^{d}\rho_{j}^{\text{agg}}$, where $d$ is the number of input features.
Step 2: XGBoost with SHAP Weighting. In this step, we enhance the RuleFit model by replacing its traditional gradient boosting component with XGBoost [36, eXtreme Gradient Boosting], an optimized implementation of gradient boosting that introduces regularization, parallel processing, and efficient handling of missing data. In particular, XGBoost allows for feature weighting during rule selection, where to each feature $c_{i}$ is possible to assign a weight, which in our case is derived from the aggregated Shapley values $\bar{\rho_{i}}$ computed in Step 1.
@@ 5 Experimental Setup
We designed an experiment to evaluate the effectiveness of global XAI methods in detecting SDG-related biases within prominent general-purpose LLMs: (i) GPT-4o, (ii) GPT-4o mini, (iii) GPT-3.5 Turbo, (iv) Llama-3.1-70B, (v) Llama-3.1. The process (see Figure 1) begins with selecting a diverse set of topics $\mathcal{T}$ related to the UN SDGs. Then, to establish a ground truth for bias detection, we introduce biases $b_{j}$ into an LLM $f$ through system instructions, creating modified models $f_{b_{j}}$. These biases vary in complexity, incorporating: (i) univariate, (ii) conjunctivemultivariate, and (iii) non-convex/disconnected (e.g., modulus operator) relationships to simulate realistic bias scenarios. Using the instructions, we then obtain explanations $\mathbf{y}_{i}=f_{b_{j}}(\mathbf{x}_{i})$, producing a corresponding set of biased responses. These inputs and outputs are further abstracted using mapping functions $\phi_{x}(\mathbf{x}_{i})$ and $\phi_{y}(\mathbf{y}_{i})$ as defined in Section 3. To mitigate stochastic effects, we set sampling temperature and top-p to 0. Next, global XAI algorithms (e.g., RuleSHAP) are applied to each biased model $f_{b_{j}}$ to detect the injected biases. Further details on the experimental setup are provided below.
@@
Topic extraction. For the topic selection phase, we focus on three domains—climate action (SDG 13), good health and well-being (SDG 3), and gender equality (SDG 5)—each corresponding to one of three distinct SDGs of the UN. These domains were chosen for their global relevance and the diverse range of topics they encompass. Within each domain, we categorize topics according to the dimensions defined in Section 3, such as whether a topic is conceptually dense, technically complicated, socially controversial, or emotionally charged, among others. Each of these dimensions is scored on a Likert scale indicating the degree to which the dimension applies to a particular topic.
Following the workaround from Section 3, we determined the required sample size for $\mathcal{T}$ through an a priori power analysis, targeting a minimum power of 0.8, a medium effect size (0.35), and $\alpha\leq 0.05$. This indicated that about 60 data points per dimension are needed. Accordingly, each LLM was tasked with generating at least n_topics = 60 unique topics per domain and score. The prompt template and hyperparameters are detailed in Appendix C. This process was repeated across the five possible scores (1–5) for each dimension to ensure comprehensive coverage. Since topics for one dimension are abstracted across others, we eventually achieve a final sample size of $60\times 11=660$ topics; well beyond the minimum required. Thus, if many similar data points (in the abstracted input space) yield outputs with the same bias-related properties, it becomes increasingly likely that their abstractions are truly correlated with these properties.
After collecting the topics, near-duplicates were removed by calculating the semantic similarity among all topics with a transformer-based model, i.e., all-MiniLM-L6-v2 [37], applying a cosine similarity threshold of 0.9 (i.e., 90% similarity). Any topic exceeding it is considered too similar and removed. This ensures that the final list (see statistics in Table 1) avoids redundancy.
@@
Injecting bias via system instructions. Let $f$ be an LLM, and let $B=\{b_{1},b_{2},b_{3}\}$ be a set of bias-inducing system instructions of increasing non-linearity. The biased variant of $f$ under instruction $b_{j}$ is denoted $f_{b_{j}}$. It consists of a Python wrapper around $f(\mathbf{x}_{i})$ that sends instruction $b_{j}$ if the input abstraction scores produced by $\phi_{x}(\tau_{i})$ satisfy the preconditions required by $b_{j}$, where $\mathbf{x}_{i}=q(\tau_{i})$. This procedure guarantees the presence of ground truth biases we can use for assessing XAI methods.
@@
Baseline XAI methods. Global rule extraction methods such as GLocalX [38], Skope-Rules [39], Bayesian Rule Sets [40], and FIGS [41] are widely used for models with categorical outputs. However, we focus on more generic, numerical outputs, where the landscape of XAI solutions is sparser. In this case, researchers typically rely on linear regression, decision tree surrogates, or other tree-based methods. Although linear regression can extract rules, it fails to capture the non-linear biases defined in Section 5, so we did not consider it as a baseline. Decision trees, instead, can handle non-linearity. Methods like MRE [42] and PALM [43] build on decision trees but do not provide implementation code, so we excluded them. RuleFit [11], instead, relies on gradient boosting and it is open source.
Hence, for our study we adopt RuleFit and decision-tree-based surrogate models as baseline explainable approaches. Furthermore, as an additional baseline, we incorporate RuleFit enhanced with XGBoost instead of traditional gradient boosting, which is equivalent to RuleSHAP with feature weights set to 1. Additionally, despite SHAP not producing explicit rule-based explanations, we decided to include it in a separate evaluation to understand to what extent its global explanations can capture the importance of features involved in bias rules. To perform an ablation study on RuleSHAP’s main components, we consider two modified versions of the RuleSHAP. The first variant omits Step 2 (cf. Section 4), excluding SHAP-driven feature weighting in XGBoost. The second variant excludes Step 3, i.e., no SHAP-Aware LASSO Regression.
Evaluation metrics. To evaluate baseline XAI methods (cf. Section 5) and RuleSHAP in identifying bias rules in LLMs, we use Mean Reciprocal Rank (MRR) [44, 45] to measure faithfulness and rule-based conciseness [46, 47] to assess compactness, which impacts interpretability. This follows standard practice in global (rule extraction) XAI [47]. MRR is used to assess how well an XAI method recovers and ranks injected bias rules. Given a set of biases $B$, let $r_{j}$ be the rank of the correct bias rule for $b_{j}\in B$ in the extracted rule set $R$. The reciprocal rank at $k\in\{1,3,10\}$ is $\mathrm{RR@}k(b_{j})=\mathbf{1}[r_{j}\leq k]/r_{j}$, and the mean reciprocal rank at top-$k$ is $\mathrm{MRR@}k\;=\;\frac{1}{|B|}\sum_{b_{j}\in B}\mathrm{RR@}k(b_{j})$. MRR measures only true positives via exact matches, underestimating but never overestimating faithfulness. Rule-based conciseness instead quantifies the number of rules extracted by an XAI method. A concise $|R|$ reduces cognitive load, with smaller $|R|$ indicating more efficient XAI [48, 49].
@@ 6 Results and Analysis
We first assess how well SHAP highlights bias-inducing features, then compare RuleSHAP to baselines on interpretability (rule compactness) and bias detection (MRR). Finally, we present an ablation study evaluating the contributions of each RuleSHAP component.
Global SHAP. We measure biased feature ranking with MRR@1, MRR@3, and MRR@10. At MRR@1, SHAP scores 0.41 on the GPTs and 0.33 on the Llamas. At MRR@3, it gets 0.48 (GPT-3.5 turbo), 0.59 (GPT-4o mini), 0.52 (GPT-4o), and $\approx$ 0.45 (Llama-3.1). MRR@10 exceeds 0.5 across all models: 0.55 (GPT-3.5 turbo), 0.62 (GPT-4o mini), 0.59 (GPT-4o), and $\approx$ 0.53 (Llama-3.1); more at Appendix F. Sorting feature importance by upper bound (mean + standard deviation) yields the highest MRRs, outperforming mean or max and motivating its use in RuleSHAP (cf. Section 4).
RuleSHAP vs. Baselines. We compared RuleSHAP against the baselines from Section 5. Table 2 reports two metrics: (i) rule-set size (smaller is better for interpretability) and (ii) MRR at different cut-offs (higher indicates better ranking). Across all LLMs, RuleSHAP consistently outperforms baselines in MRR, achieving the highest score ($\mathrm{MRR@1}=0.78$) on GPT-4o. In contrast, decision tree surrogates yield the lowest MRR, consistently scoring 0. RuleSHAP also produces more compact rule sets than RuleFit, with the smallest set (943 rules) on GPT-3.5 turbo. These results confirm RuleSHAP as both more faithful and interpretable than baselines. Statistical analysis via the Wilcoxon test shows RuleSHAP is significantly more faithful than RuleFit across RR@1, RR@3, and RR@10 (p-values: 0.003, <0.001, <0.001; moderate effect sizes 0.366–0.424), with results remaining significant after Holm-Bonferroni correction. Finally, all XAI methods detect univariate biases effectively (see Appendix G). However, baselines struggle with conjunctive and non-convex biases: RuleFit MRR@1 averages 0.4 (univariate), 0.32 (conjunctive), and 0.26 (non-convex); SHAP improves to 0.4, 0.4, and 0.36, respectively. This observation motivated the creation of RuleSHAP in the first place, which integrates SHAP with RuleFit, boosting performance to 0.66, 0.72, and 0.4.
Table 2: XAI methods comparison. The best results per LLM, column-wise, are in bold if MRR $>0$.
@@
GPT-3.5 turbo Decision Tree 987 0 0 0
@@
GPT-4o mini Decision Tree 1005 0 0 0
@@
Llama 3.1 Decision Tree 1058 0 0 0
@@
Llama 3.1 70B Decision Tree 648 0 0 0
@@
Correlation Certificates. Since the injected bias rules are non-linear, we used distance correlation [30] to compute the certificates required by the workaround in Section 3. For univariate biases, we observe medium-to-strong (0.32–0.86) correlations between topic commonality and explanation length, except for Llama-3.1 (0.15) and GPT-3.5 (0.23). For conjunctive and non-convex biases, topic positivity and subjectivity scores are strongly correlated (0.73–0.89), with explanation length showing medium-to-strong correlations (up to 0.54) with topic positivity or commonality. For non-convex biases, topic interdisciplinarity strongly correlates with readability (0.56–0.86), except for Llama-3.1 (0.05). All correlations are statistically significant ($p<0.01$) after Bonferroni correction ($\alpha=0.05$); see Appendix H for details. Additionally, Spearman analyses reveal significant ($p\ll 0.05$) correlations between LLM-generated bias estimates and proxy metrics. Framing effect strongly correlates with subjectivity ($r\approx 0.79$) and moderately negatively with sentiment ($r\approx-0.41$). Information overload moderately correlates with explanation length ($r\approx 0.60$) and the Gunning fog index ($r\approx 0.47$). Oversimplification shows moderate correlation with subjectivity ($r\approx 0.36$) and weaker correlation with Gunning fog index ($r\approx 0.23$), being the least aligned with proxy metrics.

### v2

@@ 4 RuleSHAP
SHAP [12] is a model-agnostic XAI method based on game theory. Given a model $f$ and instance $u$, it assigns each feature $u_{i}$ a Shapley value $\rho_{i}$, representing its average contribution to $f(u)$ across all feature coalitions. This is estimated by replacing features with predefined background values (not necessarily 0) to measure the impact of their removal. For global attributions, one can average SHAP’s local attributions [36], providing an overall feature-importance measure. Since exact SHAP is computationally expensive and assumes feature independence, applications often rely on approximations such as KernelSHAP (weighted linear regression) or the permutation SHAP explainer (used in our experiments), which limits permutations to second-order interactions. To address bias from violated independence assumptions, the official SHAP documentation [37] recommends grouping correlated features through hierarchical partitioning (as we do). These clusters are defined by shared target information, with joint credit assigned at the cluster level.
RuleFit [14] is a global XAI method that extracts the rules (see Appendix F) a model $f$ follows to produce outputs $\mathbf{v}$ from inputs $\mathbf{u}$. It does this by leveraging features within $\mathbf{u}$, where each input vector consists of multiple features $c_{1},c_{2},\dots,c_{m}$. RuleFit combines gradient boosting [38] with sparse linear modelling to build interpretable models. Gradient boosting iteratively constructs decision trees, where each split minimizes impurity or maximizes loss reduction. A decision path from root to leaf defines a binary rule, e.g., $(c_{i}\leq\theta_{1})\land(c_{j}>\theta_{2})\land\dots$, where $\theta_{1},\theta_{2},\dots$ are threshold values. After extracting rules, RuleFit forms a binary design matrix $\mathbf{X}$, with rows as instances and columns as rules. Each entry $X_{k,n}$ is 1 if the $k$-th rule is activated for the $n$-th instance, otherwise 0. $\mathbf{X}$ has dimensions $|R|\times|\mathbf{u}|$ (where $R$ is the set of extracted rules), capturing the presence of each rule across all instances. Then, given the coefficient vector $\mathbf{w}$ and an hyper-parameter $\alpha$ to control for sparsity, RuleFit applies a LASSO regression [39] to select a sparse set of predictive rules by solving: $\min_{\mathbf{w}}\ \frac{1}{2}\|\mathbf{v}-\mathbf{X}\mathbf{w}\|_{2}^{2}+\alpha\|\mathbf{w}\|_{1}.$
@@
Step 1: Global Shapley Values Aggregation. We first compute SHAP values for all available data points using one of the SHAP explainers available in the official repository. Specifically, for each feature $c_{i}$ and instance $\mathbf{u}_{k}$, we obtain the Shapley value $\rho_{(i,k)}$, using the minimum feature values from $\mathcal{T}$ as the background since all features are, by construction, ordered. Because we work with abstracted features, SHAP cannot directly estimate the average marginal contribution $\rho_{(i,k)}$, by simple feature “removal”, sending a new $\widehat{\mathbf{u}}_{k}$-vector to the LLM for evaluation.
Instead, following the workaround from §3, our strategy is to guarantee sufficient redundancy in $\mathcal{T}$ so that we can find multiple points $j$ in $\mathcal{T}$ for which $\|\widehat{\mathbf{u}}_{k}-\mathbf{u}_{j}\|_{2}$ is minimal ($\approx 0$) while matching the background values considered by SHAP perturbations. SHAP can then randomly select one of these points for contribution estimation, mimicking the non-determinism of LLMs and enabling a statistical assessment of how input abstractions correlate with bias-related properties. Specifically, we aggregate the $\rho_{(i,k)}$ values across all instances by calculating the mean of the absolute Shapley values, adding the standard deviation to capture the upper bound of each feature’s importance: $\rho_{i}^{\text{agg}}=\frac{1}{N}\sum_{k=1}^{N}|\rho_{(i,k)}|+\text{std}\left(\{|\rho_{(i,k)}|\}_{k=1}^{N}\right)$, where $N$ is the total number of data points $\mathbf{u}$. The aggregated Shapley values $\rho_{i}^{\text{agg}}$ are then normalized to the interval $(0,1]$ by: $\bar{\rho_{i}}=\rho_{i}^{\text{agg}}/\sum_{j=1}^{d}\rho_{j}^{\text{agg}}$, where $d$ is the number of input features.
Step 2: XGBoost with SHAP Weighting. In this step, we enhance the RuleFit model by replacing its traditional gradient boosting component with XGBoost [40, eXtreme Gradient Boosting], an optimized implementation of gradient boosting that introduces regularization, parallel processing, and efficient handling of missing data. In particular, XGBoost allows for feature weighting during rule selection, where to each feature $c_{i}$ is possible to assign a weight, which in our case is derived from the aggregated Shapley values $\bar{\rho_{i}}$ computed in Step 1.
@@ 5 Experimental Setup
We evaluate the effectiveness of global XAI in detecting belief-related heuristics across five popular general-purpose LLMs (GPT-4o, GPT-4o mini, GPT-3.5-Turbo, Llama-3.1-70B, Llama-3.1). As shown in Fig. 1, we first sample a large set of SDG-related topics $\mathcal{T}$3). To obtain ground truth, we hard-code bias-inducing rules $b_{j}$ into a base LLM $f$ via system instructions, producing modified models $f_{b_{j}}$. These rules vary in complexity, incorporating nonlinear relationships of three different types: univariate, conjunctive multivariate, and non-convex/disconnected (e.g., modulus operator). Using $b_{j}$, we generate explanations $\mathbf{y}_{i}\!=\!f_{b_{j}}(\mathbf{x}_{i})$ and abstract inputs/outputs via $\phi_{x}(\mathbf{x}_{i})$ and $\phi_{y}(\mathbf{y}_{i})$. Finally, we apply global XAI (e.g., RuleSHAP) to each biased model $f_{b_{j}}$ to detect injected heuristics by analysing LLM behaviour across all topics $\mathcal{T}$. Further experimental setup details follow.
@@
Topic extraction. For the topic selection phase, we focus on three domains—climate action (SDG 13), good health and well-being (SDG 3), and gender equality (SDG 5)—each corresponding to one of three distinct SDGs of the UN. These domains were chosen for their global relevance and the diverse range of topics they encompass. Within each domain, we categorize topics according to the dimensions defined in §3, such as whether a topic is conceptually dense, technically complicated, socially controversial, or emotionally charged, among others. Each of these dimensions is scored on a Likert scale indicating the degree to which the dimension applies to a particular topic.
Following the workaround in §3, we determined the required sample size for $\mathcal{T}$ through an a priori power analysis, targeting power of .8, a medium effect size (.35.35), and $\alpha\!\leq\!.05$. This indicated that about 60 data points per dimension are needed to ensure sufficient power for correlation certificates. Accordingly, each LLM was tasked with generating at least 60 unique topics per domain and score.
The prompt template and hyperparameters are detailed in Appendix C. This process was repeated across the five possible scores (1–5) for each dimension to ensure comprehensive coverage. Since topics for one dimension are abstracted across others, we eventually achieve a final sample size of $60\times 11=660$ topics; well beyond the minimum required. Thus, if many similar data points (in the abstracted input space) yield outputs with the same bias-related properties, it becomes increasingly likely that their abstractions are truly correlated with these properties.
After collecting the topics, near-duplicates were removed by calculating the semantic similarity among all topics with a transformer-based model, i.e., all-MiniLM-L6-v2 [41], applying a cosine similarity threshold of 90%. Any topic exceeding it is considered too similar and removed. This ensures that the final list (see statistics in Table 1) avoids redundancy.
@@
Injecting bias via system instructions. Let $B=\{b_{1},b_{2},b_{3}\}$ be a set of bias-inducing system instructions of increasing non-linearity. The biased variant of the LLM $f$ under instruction $b_{j}$ is denoted $f_{b_{j}}$. It consists of a Python wrapper around $f(\mathbf{x}_{i})$ sending $b_{j}$ if the input abstraction scores produced by $\phi_{x}(\tau_{i})$ satisfy the preconditions required by $b_{j}$ (see Appendix D), where $\mathbf{x}_{i}=q(\tau_{i})$. This procedure guarantees the presence of ground-truth biases we can use for assessing XAI methods.
@@
We decode explanations and abstraction scores deterministically (T=0T\!=\!0, top-p =0\!=\!0), so outputs reflect instructions and LLM parameters (i.e., beliefs), not randomness. Higher temperatures (T=2T\!=\!2) are nonetheless possible but induce off-instruction drift (especially in smaller models like Llama-3.1), weakening correlation certificates; rules with no correlation certificate are red-flagged (§3 and §6).
Baseline XAI methods. Global rule extraction methods such as GLocalX [42], Skope-Rules [43], Bayesian Rule Sets [44], and FIGS [45] are widely used for models with categorical outputs. However, we focus on more generic, numerical outputs, where the landscape of XAI solutions is sparser. In this case, researchers typically rely on linear regression or other tree-based methods (e.g., decision trees). Although linear regression can extract rules, it fails to capture the non-linear biases defined in §5, thus we did not consider it as a baseline. Decision trees, instead, can handle non-linearity. Methods like MRE [46] and PALM [47] use them but lack implementation code, so are excluded.
Hence, for our study we adopt RuleFit [14] (which relies on gradient boosting and it is open source) and decision-tree-based surrogate models as baselines. Moreover, as an additional baseline, we incorporate RuleFit enhanced with XGBoost instead of traditional gradient boosting, which is equivalent to RuleSHAP with feature weights set to 1. Despite SHAP not producing explicit rule-based explanations, we also decided to include it in a separate evaluation to understand to what extent its global explanations can capture the importance of features involved in bias rules. To perform an ablation study on RuleSHAP’s main components, we consider two modified versions of the RuleSHAP. The first variant omits Step 2 (§4), excluding SHAP-driven feature weighting in XGBoost. The second variant excludes Step 3, i.e., no SHAP-Aware LASSO.
Evaluation metrics. To evaluate baseline XAI methods (§5) and RuleSHAP in identifying bias rules in LLMs, we use Mean Reciprocal Rank (MRR) [48, 49] to measure faithfulness and rule-based conciseness [50, 51] to assess compactness, which impacts interpretability. This follows standard practice in global (rule extraction) XAI [51]. MRR is used to assess how well an XAI method recovers and ranks injected bias rules. Given a set of biases $B$, let $r_{j}$ be the rank of the correct bias rule for $b_{j}\in B$ in the extracted rule set $R$. The reciprocal rank at $k\!\in\!\{1,3,10\}$ is $\mathrm{RR@}k(b_{j})=\mathbf{1}[r_{j}\leq k]/r_{j}$, and the mean reciprocal rank at top-$k$ is $\mathrm{MRR@}k\!=\!\frac{1}{|B|}\sum_{b_{j}\in B}\mathrm{RR@}k(b_{j})$. MRR measures only true positives via exact matches, underestimating but never overestimating faithfulness. Rule-based conciseness instead quantifies the number of rules extracted by an XAI method. A concise $|R|$ reduces cognitive load, with smaller $|R|$ indicating more efficient XAI [52].
@@ 6 Results and Analysis
We first assess how well SHAP highlights bias-inducing features, then compare RuleSHAP to baselines on interpretability (rule compactness) and bias detection (MRR).
Global SHAP. We measure biased feature ranking with MRR@1, MRR@3, and MRR@10. At MRR@1, SHAP scores 0.41 on the GPTs and 0.33 on the Llamas. At MRR@3, it gets 0.48 (GPT-3.5 turbo), 0.59 (GPT-4o mini), 0.52 (GPT-4o), and 0.45 (Llama-3.1). MRR@10 exceeds 0.5 across all models: 0.55 (GPT-3.5 turbo), 0.62 (GPT-4o mini), 0.59 (GPT-4o), and 0.53 (Llama-3.1); more at Appendix G. Sorting feature importance by upper bound (mean + standard deviation) yields the highest MRRs, empirically outperforming mean or max and motivating its use in RuleSHAP4).
@@
GPT-3.5
turbo
Decision Tree 987 0 0 0
@@
GPT-4o
mini
Decision Tree 1005 0 0 0
@@
Llama
3.1
Decision Tree 1058 0 0 0
@@
Llama
3.1 70B
Decision Tree 648 0 0 0
@@
Table 2: XAI methods comparison. The best results per LLM, column-wise, are in bold if MRR >0>0.
RuleSHAP vs. Baselines. We compared RuleSHAP against the baselines from §5. Table 2 reports two metrics: (i) rule-set size (smaller is better for interpretability) and (ii) MRR at different cut-offs (higher indicates better ranking). Across all LLMs, RuleSHAP consistently outperforms baselines in MRR, achieving the highest score ($\mathrm{MRR@1}=0.78$) on GPT-4o. In contrast, decision tree surrogates yield the lowest MRR, consistently scoring 0. RuleSHAP also produces more compact rule sets than RuleFit , with the smallest set (943 rules) on GPT-3.5 turbo. These results confirm RuleSHAP as both more faithful and interpretable than baselines. Statistical analysis via the Wilcoxon test shows RuleSHAP is significantly more faithful than RuleFit across RR@1, RR@3, and RR@10 (p-values: 0.003, <0.001, <0.001; moderate effect sizes 0.366–0.424), with results remaining significant after Holm-Bonferroni correction. Finally, all XAI methods detect univariate biases effectively (see Appendix H). However, baselines struggle with conjunctive and non-convex biases: RuleFit MRR@1 averages 0.4 (univariate), 0.32 (conjunctive), and 0.26 (non-convex); SHAP improves to 0.4, 0.4, and 0.36, respectively. This observation motivated the creation of RuleSHAP in the first place, which integrates SHAP with RuleFit, boosting performance to 0.66, 0.72, and 0.4.
Correlation Certificates. Since the injected bias rules are non-linear, we used distance correlation [29] to compute the certificates required by the workaround in §3. For univariate biases, we observe medium-to-strong (0.32–0.86) correlations between topic commonality and explanation length, except for Llama-3.1 (0.15) and GPT-3.5 (0.23). For conjunctive and non-convex biases, topic positivity and subjectivity scores are strongly correlated (0.73–0.89), with explanation length showing medium-to-strong correlations (up to 0.54) with topic positivity or commonality.
For non-convex biases, topic interdisciplinarity strongly correlates with readability (0.56–0.86), except for Llama-3.1 (0.05). All correlations are statistically significant (p<0.01p\!<\!0.01) after Bonferroni correction ($\alpha\!=\!0.05$); see Appendix I for details. Additionally, Spearman analyses reveal significant ($p\!\ll\!0.05$) correlations between LLM-generated bias estimates and proxy metrics. Framing effect strongly correlates with subjectivity (r=0.79r\!=\!0.79) and moderately negatively with sentiment (r=0.41r\!=\!-0.41). Information overload moderately correlates with explanation length (r=0.60r\!=\!0.60) and the Gunning fog index (r=0.47r\!=\!0.47). Oversimplification shows moderate correlation with subjectivity (r=0.36r\!=\!0.36) and weaker correlation with Gunning fog index (r=0.23r\!=\!0.23), being the least aligned with proxy metrics.